OpenStudy (anonymous):

Use the Change of Base Formula to approximate each of the following to two decimal places. 1) log2 6 = 2) log2 13 = I have a whole worksheet of these, anyone remember them, can u show me how to solve them>

OpenStudy (anonymous):

@primeralph I used my friends account

OpenStudy (anonymous):

\[b^x=y\iff x=\frac{\ln(y)}{\ln(b)}\]

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

How are you defining solve? If you just want an approximate solution, to what accuracy?

OpenStudy (anonymous):

\[\log_b(A)=\frac{\ln(A)}{\ln(b)}\]

OpenStudy (anonymous):

two decimal places

OpenStudy (anonymous):

so for example \[\log_2(6)=\frac{\ln(6)}{\ln(2)}\]

OpenStudy (anonymous):

@satellite73 so it would be 3?

OpenStudy (primeralph):